hhanan:

Suppose the population of a small town is 567 in 2011. The population decreases at a rate of 1.5% every year. What will be the population of the town in 2020? Round your answer to the nearest whole number.

ScamTheMan:

This question is exponential, and although we could do it the incredibly long way of multiplying 0.015 by 567 9 times, we can set up an equation. Since it's exponential, we can already set up an equation. Our formula for setting up an equation is as follows: \[f(x)=a(1+r)^x\] a = Initial Amount r = Growth Rate x = Time Intervals Now we can set up our equation by filing in the values: \[f(x)=567(1-0.015)^9\] Now we can just put this into an calculator and get an answer: \[f(x)=567(0.985)^9\] \[f(x)=494.8\] Meaning our answer is 494.8, which we can easily round up to 495

ScamTheMan:

I felt like this explanation may be a bit confusing for some people, so if you need clarification, let me know